# Gradient descent induces alignment between weights and the empirical NTK for deep non-linear networks

### v1

@@ Abstract
Understanding the mechanisms through which neural networks extract statistics from input-label pairs is one of the most important unsolved problems in supervised learning. Prior works have identified that the gram matrices of the weights in trained neural networks of general architectures are proportional to the average gradient outer product of the model, in a statement known as the Neural Feature Ansatz (NFA). However, the reason these quantities become correlated during training is poorly understood. In this work, we explain the emergence of this correlation. We identify that the NFA is equivalent to alignment between the left singular structure of the weight matrices and a significant component of the empirical neural tangent kernels associated with those weights. We establish that the NFA introduced in prior works is driven by a centered NFA that isolates this alignment. We show that the speed of NFA development can be predicted analytically at early training times in terms of simple statistics of the inputs and labels. Finally, we introduce a simple intervention to increase NFA correlation at any given layer, which dramatically improves the quality of features learned.
@@ 1 Introduction
Neural networks have emerged as the state-of-the-art machine learning methods for seemingly complex tasks, such as language generation (Brown et al., 2020), image classification (Krizhevsky et al., 2012), and visual rendering (Mildenhall et al., 2021). The precise reasons why neural networks generalize well have been the subject of intensive exploration, beginning with the observation that standard generalization bounds from statistical learning theory fall short of explaining their performance (Zhang et al., 2021).
@@
A promising line of work emerged in the form of the neural tangent kernel (NTK), connecting neural networks to kernels in the wide limit (Jacot et al., 2018; Chizat et al., 2019). However, subsequent research showed that the success of neural networks relies critically on aspects of learning which are absent in kernel approximations (Ghorbani et al., 2019; Allen-Zhu and Li, 2019; Yehudai and Shamir, 2019; Li et al., 2020; Refinetti et al., 2021). Other work showed that low width suffices for gradient descent to achieve arbitrarily small test error (Ji and Zhu, 2020), further refuting the idea that extremely wide networks are necessary.
@@
Subsequently, the success of neural networks has been largely attributed to feature learning - the ability of neural networks to learn statistics, measurements, and representations of data which are useful for downstream tasks. However, the specific mechanism through which features are learned is an important unsolved problem in deep learning theory. A number of works have studied the abilities of neural networks to learn features in structured settings (Abbe et al., 2022; Ba et al., 2022; Nichani et al., 2023; Barak et al., 2022; Damian et al., 2022; Moniri et al., 2023; Parkinson et al., 2023). Some of that work proves strict separation in terms of sample complexity between neural networks trained with stochastic gradient descent and kernels (Mousavi-Hosseini et al., 2022).
@@
The work above studies simple structure, such as learning from low-rank data or functions that are hierarchical compositions of simple elements. Recent work makes a big step towards generalizing these assumptions by proposing the neural feature ansatz (NFA) (Radhakrishnan et al., 2022; Beaglehole et al., 2023), a general structure that emerges in the weights of trained neural networks. The NFA states that the gram matrix of the weights at a given layer (known as the neural feature matrix (NFM)) is aligned with the average gradient outer product (AGOP) of the network with respect to the input to that layer. In particular, the NFM and AGOP are highly correlated in all layers of trained neural networks of general architectures, including practical models such as AlexNet (Krizhevsky et al., 2012) and VGG (Simonyan and Zisserman, 2014).
@@
A major missing element of this theory is the reason the AGOP and NFM become correlated during training. In this paper, we explain the emergence of this correlation. We establish that the NFA is equivalent to alignment between the left singular structure of the weight matrices and the pre-activation neural tangent kernel (PTK), which is closely related to empirical neural tangent kernel (ENTK) (Section 2). By introducing a modified NFA, we demonstrate that the original NFA is driven by alignment of the left singular vectors of the weight matrices to with the PTK (Section 2). Further, we show that the alignment speed can be understood analytically in terms of the statistics of the data and labels at early times, and can be manipulated theoretically (Section 3).
@@
Finally, we force the strength of the NFA alignment to improve using two interventions: small initialization and Speed Limited Optimization, a layerwise gradient normalization scheme (Section 4). This suggests a path towards rational design of architectures and training procedures that maximize this notion of feature learning.
2 Alignment between the weight matrices and the empirical NTK
@@ 2.1 Preliminaries
We consider fully-connected neural networks with a single output of depth $L\geq 1$, where $L$ is the number of hidden layers, written $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. We write the input to layer $\ell\in\{0,\ldots,L-1\}$ as $x_{\ell}$, where $x_{0}\equiv x$ is the original datapoint, and the pre-activation as $h_{\ell}(x)$. Then,
$h_{\ell}(x)=W^{(\ell)}x_{\ell},\qquad x_{\ell+1}=\phi(h_{\ell}(x))~{},$
where $\phi$ is an element-wise nonlinearity, $W^{(\ell)}\in\mathbb{R}^{k_{\ell+1}\times k_{\ell}}$ is a weight matrix, and $k_{\ell}$ is the hidden dimension at layer $\ell$. We restrict $k_{L+1}$ to be the number of output logits, and set $k_{0}=d$, where $d$ is the input dimension of the data. Note that $f(x)=h_{L+1}(x)$. We train $f$ by gradient descent on a loss function $\mathcal{L}(\theta,X)$, where $X$ is an input dataset, and $\theta$ is the collection of weights.
We consider a supervised learning setup where we are provided $n$ input-label pairs $(x^{(1)},y^{(1)}),\ldots,(x^{(n)},y^{(n)})\in\mathbb{R}^{d}\times\mathbb{R}$. We denote the inputs $X\in\mathbb{R}^{n\times d}$ and the labels $y\in\mathbb{R}^{n\times 1}$. We train a fully-connected neural network to learn the mapping from inputs to labels by minimizing a standard loss function, such as mean-squared-error or cross-entropy, on the dataset.
@@
One can define two objects associated with neural networks that capture learned structure. For a given layer $\ell$, the neural feature matrix (NFM) $F_{\ell}$ is the gram matrix of the columns of the weight matrix $W^{(\ell)}$, i.e. $F_{\ell}\equiv(W^{(\ell)})^{\top}W^{(\ell)}$. The second fundamental object we consider is the average gradient outer product (AGOP) $\bar{G}_{\ell}$, defined as $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}^{\top}$. To understand the structure of these objects, consider the following chain-monomial low-rank task:
@@
where the data inputs are sampled from an isotropic Gaussian distribution $\mu=\mathcal{N}(0,I)$. In this case, both objects capture the coordinates on which the target function depends (Figure 1). Prior work has shown that in trained neural networks, these objects will be proportional to each other to varying degree of approximation. This notion is formalized in the Neural Feature Ansatz (NFA):
Ansatz 1 (Neural Feature Ansatz (Radhakrishnan et al., 2022)).
The Neural Feature Ansatz states that, for all layers $\ell\in[L]$ of a fully-connected neural network with $L$ hidden layers trained on input data $x^{(1)},\ldots,x^{(n)}$,
$(W^{(\ell)})^{\top}W^{(\ell)}\propto\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}^{\top}~{},$
The left and right hand sides of this equation are the NFM and AGOP with respect to the input of layer $\ell$. Here, $\frac{\partial f(x)}{\partial x_{\ell}}\in\mathbb{R}^{k_{\ell}\times 1}$ denotes the gradient of the function $f$ with respect to the intermediate representation $x_{\ell}$. For simplicity, we may concatenate these gradients into a single matrix $\frac{\partial f(X)}{\partial x_{\ell}}\in\mathbb{R}^{n\times k_{\ell}}$. Note we consider scalar outputs in this work, though the NFA relation is identical when there are $c\geq 1$ outputs, where in this case $\frac{\partial f(x)}{\partial x_{\ell}}\in\mathbb{R}^{k_{\ell}\times c}$ is the input-output Jacobian of the model $f$.
The NFA is equivalent to the statement that, for every layer $\ell$, the NFM $F_{\ell}$ and AGOP $\bar{G}_{\ell}$, will have cosine similarity, or correlation,
$\rho\left(F_{\ell},\bar{G}_{\ell}\right)=1~{},$
where the correlation $\rho$, with range $[-1,1]$, is defined as
$\rho\left(A,B\right)=\text{tr}\left(A^{\top}B\right)\cdot\text{tr}\left(A^{\top}A\right)^{-1/2}\cdot\text{tr}\left(B^{\top}B\right)^{-1/2}~{}.$
for any two matrices $A,B\in\mathbb{R}^{d_{1}\times d_{2}}$, for any $d_{1}$ and $d_{2}$.
We note that while the NFA states exact proportionality, in practice, the NFM and AGOP are highly correlated with correlation less than $1$ (see e.g. Figure 2), where the final correlation depends on many aspects of training and architecture choice. For example, the final value of the UC-NFA can be sensitive to the magnitude of the initial weights (e.g. first column of Figure 2). We parameterize this magnitude by the initialization scale, $s_{\ell}$, for layer $\ell$, where the initial weights are sampled i.i.d. as $W^{(\ell)}_{ij}\sim\mathcal{N}\left(0,\frac{s_{\ell}}{k_{\ell}}\right)$.
The relation between the NFM and the AGOP is significant, in part, because for a neural network that has learned enough information about the target function, the AGOP of this model with respect to the first-layer inputs will approximate the expected gradient outer product (EGOP) of the target function (Yuan et al., 2023). In particular, as we will see later with the example of a low rank polynomial, the EGOP of the target function contains task-specific structure that is completely independent of the model used to estimate it. Where the labels are generated from a particular target function $y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ on data sampled from a distribution $\mu$, the EGOP is equal to,
If the NFA holds, the correlation of the EGOP and the AGOP at the end of training also implies high correlation between the NFM of the first layer and the EGOP, so that the NFM has encoded this task-specific structure.
@@
To demonstrate the significance of the approximate proportionality between the NFM and the AGOP in successfully trained networks, we return to the chain-monomial low rank task. In this case, $\operatorname{EGOP}(y,\mu)$ will be rank $r$, where $r$ is much less than the ambient dimension (Figure 1). Similarly, the AGOP of the trained model resembles the EGOP. In this case the NFA implies that the NFM also resembles the EGOP. Therefore, the neural network has learned the model-independent and task-specific structure of the chain-monomial task in the right singular values and vectors of the first layer weight matrix, as these are determined by the NFM. As a demonstration of this fact, the NFM of a trained neural network can be recovered up to high correlation from the AGOP of a fixed kernel method on real datasets (Radhakrishnan et al., 2022)
@@
This insight has inspired iterative kernel methods which can match the performance of fully-connected networks (Radhakrishnan et al., 2022, 2024) and can close the gap with convolutional networks (Beaglehole et al., 2023). Additional prior works demonstrate the benefit of including the AGOP features to improve feature-less predictors (Hristache et al., 2001; Trivedi et al., 2014; Kpotufe et al., 2016). Additionally, because the NFM is correlated with the AGOP, the AGOP can be used to recover the features from feature-less methods, such as kernel machines.
@@ 2.2 Alignment decomposition
In order to understand the NFA, it is useful to decompose the AGOP. Doing so will allow us to show that the NFA can be interpreted as an alignment between weight matrices and the pre-activation tangent kernel (PTK), which is closely related to the neural tangent kernel (NTK).
For any layer $\ell$, we can re-write the AGOP as,
$\bar{G}_{\ell}=(W^{(\ell)})^{\top}K^{(\ell)}W^{(\ell)},~{}K^{(\ell)}\equiv\frac{\partial f(X)}{\partial h_{\ell}}^{\top}\frac{\partial f(X)}{\partial h_{\ell}}$
@@
Proposition 2 (Alignment decomposition of NFA).

### v2

# Feature learning as alignment: a structural property of gradient descent in non-linear neural networks

@@ Abstract
Understanding the mechanisms through which neural networks extract statistics from input-label pairs through feature learning is one of the most important unsolved problems in supervised learning. Prior works demonstrated that the gram matrices of the weights (the neural feature matrices, NFM) and the average gradient outer products (AGOP) become correlated during training, in a statement known as the neural feature ansatz (NFA). Through the NFA, the authors introduce mapping with the AGOP as a general mechanism for neural feature learning. However, these works do not provide a theoretical explanation for this correlation or its origins. In this work, we further clarify the nature of this correlation, and explain its emergence. We show that this correlation is equivalent to alignment between the left singular structure of the weight matrices and the newly defined pre-activation tangent features at each layer. We further establish that the alignment is driven by the interaction of weight changes induced by SGD with the pre-activation features, and analyze the resulting dynamics analytically at early times in terms of simple statistics of the inputs and labels. Finally, motivated by the observation that the NFA is driven by this centered correlation, we introduce a simple optimization rule that dramatically increases the NFA correlations at any given layer and improves the quality of features learned.
**footnotetext: Correspondence to: Daniel Beaglehole (dbeaglehole@ucsd.edu)$\dagger$\$$\dagger$\$footnotetext: Work partially done as an intern at Google DeepMind.
@@ 1 Introduction
Neural networks have emerged as the state-of-the-art machine learning methods for seemingly complex tasks, such as language generation [1], image classification [2], and visual rendering [3]. The precise reasons why neural networks generalize well have been the subject of intensive exploration, beginning with the observation that standard generalization bounds from statistical learning theory fall short of explaining their performance [4].
@@
A promising line of work emerged in the form of the neural tangent kernel, connecting neural networks to kernels in the wide limit [5, 6]. However, subsequent research showed that the success of neural networks relies critically on aspects of learning which are absent in kernel approximations [7, 8, 9, 10, 11]. Other work showed that low width suffices for gradient descent to achieve arbitrarily small test error [12], further refuting the idea that extremely wide networks are necessary.
@@
Subsequently, the success of neural networks has been largely attributed to feature learning - the ability of neural networks to learn statistics, measurements, and representations of data which are useful for downstream tasks. However, the specific mechanism through which features are learned is an important unsolved problem in deep learning theory. A number of works have studied the abilities of neural networks to learn features in structured settings [13, 14, 15, 16, 17, 18, 19]. Some of that work proves strict separation in terms of sample complexity between neural networks trained with stochastic gradient descent and kernels [20].
@@
The work above studies simple structure, such as learning from low-rank data or functions that are hierarchical compositions of simple elements. Recent work makes a big step towards generalizing these assumptions by proposing the neural feature ansatz (NFA) [21, 22], a general structure that emerges in the weights of trained neural networks. The NFA states that the gram matrix of the weights at a given layer (known as the neural feature matrix or NFM) is aligned with the average gradient outer product (AGOP) of the network with respect to the input to that layer. In particular, the NFM and AGOP are highly correlated in all layers of trained neural networks of general architectures, including practical models such as VGG [23], vision transformers [24], and GPT-family models [1].
@@
A major missing element is an explanation for how and why the AGOP and NFM become correlated through training with gradient descent. In this paper, we precisely describe the emergence of this correlation. We establish that the NFA is equivalent to alignment between the left singular structure of the weight matrices and the uncentered covariance of the pre-activation tangent kernel (PTK) (Section 2) features. We then introduce the centered neural feature correlation (C-NFC) which isolates this alignment process. We show empirically that the C-NFC is close to its maximum value of $1$ at early times, and fully captures the NFA at late times for a variety of architectures (fully-connected, convolutional, and attention layers) over a diverse collection of datasets (Section 3). Our experiments suggest that the C-NFC drives the development of the NFA. Through this centering, we identify analytically how the NFA emerges as a structural property of gradient descent. We provide further evidence by showing that the dynamics of the C-NFC can be understood analytically in terms of the statistics of the data and labels at early times (Section 3), and can be manipulated theoretically (Section 4).
@@
Finally, motivated by our theory, we design an intervention to increase the influence of the C-NFC and make the NFA more robust: Speed Limited Optimization, a layerwise gradient normalization scheme (Section 5). The effectiveness of the latter update rule suggests a path towards rational design of architectures and training procedures that maximize the NFA notion of feature learning by promoting alignment dynamics.
2 Alignment between the weight matrices and the pre-activation tangents
In this section, we decompose the AGOP into the weight matrices and the feature covariance of the pre-activation tangent kernel (PTK), and demonstrate that the NFA is equivalent to alignment between the weights and these PTK features. We include a glossary of terms for ease of reference in Appendix A.
@@ 2.1 Preliminaries
We consider fully-connected neural networks with a single output of depth $L\geq 1$, where $L$ is the number of hidden layers, written $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. We write the input to layer $\ell\in\{0,\ldots,L\}$ as $x_{\ell}$, where $x_{0}\equiv x$ is the original datapoint, and the pre-activation as $h_{\ell}(x)$. Then,
$h_{\ell}(x)=W^{(\ell)}x_{\ell},\qquad x_{\ell+1}=\phi(h_{\ell}(x))~{},$ (1)
where $\phi$ is an element-wise nonlinearity, $W^{(\ell)}\in\mathbb{R}^{k_{\ell+1}\times k_{\ell}}$ is a weight matrix, and $k_{\ell}$ is the hidden dimension at layer $\ell$. We restrict $k_{L+1}$ to be the number of output logits, and set $k_{0}=d$, where $d$ is the input dimension of the data. Note that $f(x)=h_{L}(x)$ and we ignore $x_{L+1}$. We train $f$ by gradient descent on a loss function $\mathcal{L}(\theta,X)$, where $X$ is an input dataset, and $\theta$ is the collection of weights.
We consider a supervised learning setup where we are provided $n$ input-label pairs $(x^{(1)},y^{(1)}),\ldots,(x^{(n)},y^{(n)})\in\mathbb{R}^{d}\times\mathbb{R}$. We denote the network inputs (datapoints) $X\in\mathbb{R}^{n\times d}$ and the labels $y\in\mathbb{R}^{n\times 1}$. For a given network, the inputs to intermediate layers $\ell\in\{0,\ldots,L\}$ are written $X_{\ell}\in\mathbb{R}^{n\times k_{\ell}}$, where $X_{0}\equiv X$. We train a fully-connected neural network to learn the mapping from network inputs to labels by minimizing a standard loss function, such as mean-squared-error or cross-entropy, on the dataset.
@@
One can define two objects associated with neural networks that capture learned structure. For a given layer $\ell$, the neural feature matrix (NFM) $F_{\ell}$ is the gram matrix of the columns of the weight matrix $W^{(\ell)}$, i.e. $F_{\ell}\equiv(W^{(\ell)})^{\top}W^{(\ell)}$. $F_{\ell}$ depends on the right singular vectors (and corresponding singular values) of $W^{(\ell)}$. The second fundamental object we consider is the average gradient outer product (AGOP) $\bar{G}_{\ell}$, defined as $\bar{G}_{\ell}\equiv\frac{1}{n}\sum_{\alpha=1}^{n}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}\frac{\partial f(x^{(\alpha)}_{\ell})}{\partial x_{\ell}}^{\top}$.
Since both these matrices have the same dimensions ($k_{\ell+1}\times k_{\ell+1}$), we can consider their cosine similarity. We define the neural feature correlation (NFC) by:
$\rho\left(F_{\ell},\bar{G}_{\ell}\right)\equiv\text{tr}\left(F_{\ell}^{\top}\bar{G}_{\ell}\right)\cdot\text{tr}\left(F_{\ell}^{\top}F_{\ell}\right)^{-1/2}\cdot\text{tr}\left(\bar{G}_{\ell}^{\top}\bar{G}_{\ell}\right)^{-1/2}~{}.$ (NFC)
This takes on values in $[0,1]$ since $F_{\ell}$ and $\bar{G}_{\ell}$ are both PSD.
Prior work has shown that in trained neural networks, the NFC will generally be close to $1$ to varying degree of approximation. This notion is formalized in the Neural Feature Ansatz (NFA):
Ansatz 1 (Neural Feature Ansatz [21]).
The Neural Feature Ansatz states that, for all layers $\ell\in[L]$ of a fully-connected neural network with $L$ hidden layers trained on input data $x^{(1)},\ldots,x^{(n)}$, the NFC will satisfy $\rho\left(\bar{G}_{\ell},F_{\ell}\right)\approx 1~{}.$
The inputs to the covariance are the NFM and AGOP with respect to the input of layer $\ell$. Here, $\frac{\partial f(x)}{\partial x_{\ell}}\in\mathbb{R}^{k_{\ell}\times 1}$ denotes the gradient of the function $f$ with respect to the intermediate representation $x_{\ell}$. For simplicity, we may concatenate these gradients across the input dataset into a single matrix $\frac{\partial f(X)}{\partial x_{\ell}}\in\mathbb{R}^{n\times k_{\ell}}$. Note we consider scalar outputs in this work, though the NFA relation is identical when there are $c\geq 1$ outputs, where in this case $\frac{\partial f(x)}{\partial x_{\ell}}\in\mathbb{R}^{k_{\ell}\times c}$ is the input-output Jacobian of the model $f$.
We note that while the ansatz states exact proportionality, in practice, the NFM and AGOP are highly correlated with correlation less than $1$, where the final correlation depends on many aspects of training and architecture choice (discussed further in Section 3).
The relation between the NFM and the AGOP is significant, in part, because for a neural network that has learned enough information about the target function, the AGOP of this model with respect to the first-layer inputs will approximate the expected gradient outer product (EGOP) of the target function [25]. In particular, the EGOP of the target function contains task-specific structure that is completely independent of the model used to estimate it. Where the labels are generated from a particular target function $y(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$ on data sampled from a distribution $\mu$, the EGOP is defined as
If the neural feature ansatz (Ansatz 1) holds, the correlation of the EGOP and the AGOP at the end of training also implies high correlation between the NFM of the first layer and the EGOP, so that the NFM has encoded this task-specific structure.
To demonstrate the significance of high correlation between the NFM and the AGOP in successfully trained networks, we consider the following chain-monomial low-rank task:
@@
where the data inputs are sampled from an isotropic Gaussian distribution $\mu=\mathcal{N}(0,I)$. In this case, the entries $\operatorname{EGOP}(y,\mu)$ will be 00 for rows and columns outside of the $r\times r$ sub-matrix corresponding to $x_{1},\ldots,x_{r}$ (Figure 1), as $y$ does not vary with coordinates $x_{r+1},\ldots,x_{d}$. Within this sub-matrix, the diagonal entries will have value $2$, while the off-diagonal entries will be either $1$ or 00. Therefore, $\operatorname{EGOP}(y,\mu)$ will be rank $r$, where $r$ is much less than the ambient dimension.
@@
We verify for this task that the AGOP of the trained model resembles the EGOP (first and third panels of Figure 1). Here the NFA holds and the NFM (second panel) resembles the AGOP and therefore the EGOP as well. Therefore, the neural network has learned the model-independent and task-specific structure of the chain-monomial task in the right singular values and vectors of the first layer weight matrix, as these are determined by the NFM. In fact, previous works have demonstrated that the NFM of the first layer of a well-trained neural network is highly correlated with the AGOP of a fixed kernel method trained on the same dataset [21].
@@
This insight has inspired iterative kernel methods which can match the performance of fully-connected networks [21, 26, 27] and improve over fixed convolutional kernels [22]. Additional prior works demonstrate the benefit of including the AGOP features to improve feature-less predictors [28, 29, 30]. Additionally, because the NFM is correlated with the AGOP, the AGOP can be used to recover the features from feature-less methods, such as kernel machines. The AGOP has also been show to capture surprising phenomena of neural networks beyond low-rank feature learning including deep neural collapse [31].
@@ 2.2 Alignment decomposition
In order to understand Ansatz 1, it is useful to decompose the AGOP. Doing so will allow us to show that the neural feature correlation (NFC) can be interpreted as an alignment metric between weight matrices and the pre-activation tangent kernel (PTK). The PTK $\mathcal{K}^{(\ell)}(x,z)$ is defined with respect to a layer $\ell$ of a neural network and two inputs $x,z$. The kernel evaluates to:
$\mathcal{K}^{(\ell)}(x,z)\equiv\frac{\partial f(x)}{\partial h_{\ell}}\cdot\frac{\partial f(z)}{\partial h_{\ell}}~{}.$ (PTK)
When the arguments $x$ and $z$ are omitted, $\mathcal{K}^{(\ell)}\in\mathbb{R}^{n\times n}$ consists of the matrix of kernel evaluations on all pairs of datapoints. We may also consider the covariance $K^{(\ell)}\in\mathbb{R}^{k_{\ell}\times k_{\ell}}$ of the features associated with the PTK, $\frac{\partial f(X)}{\partial h_{\ell}}\in\mathbb{R}^{n\times k_{\ell}}$. In particular, we define the feature covariance of the PTK as,
$K^{(\ell)}\equiv\frac{\partial f(X)}{\partial h_{\ell}}^{\top}\frac{\partial f(X)}{\partial h_{\ell}}~{}.$ (PTK feature covariance)
Crucially, for any layer $\ell$, we can re-write the AGOP in terms of this feature covariance as,
$\bar{G}_{\ell}=(W^{(\ell)})^{\top}K^{(\ell)}W^{(\ell)}~{}.$
@@
Proposition 2 (Alignment decomposition of NFC).